# The Hitchhiker’s Guide to Program Analysis, Part II: Deep Thoughts by LLMs

### v2

@@ Abstract.
Static analysis is a cornerstone for software vulnerability detection, yet it often struggles with the classic precision-scalability trade-off. In practice, such tools often produce high false positive rates, particularly in large codebases like the Linux kernel. This imprecision can arise from simplified vulnerability modeling and over-approximation of path and data constraints. While Large Language Models (LLMs) show promise in code understanding, their naive application to program analysis yields unreliable results due to inherent reasoning limitations. We introduce BugLens, a post-refinement framework that significantly improves static analysis precision. BugLens guides an LLM to follow traditional analysis steps by assessing buggy code patterns for security impact and validating the constraints associated with static warnings. Evaluated on real-world Linux kernel bugs, BugLens raises precision from 0.10 (raw) and 0.50 (semi-automated refinement) to 0.72, substantially reducing false positives and revealing four previously unreported vulnerabilities. Our results suggest that a structured LLM-based workflow can meaningfully enhance the effectiveness of static analysis tools.
conference: ; Apr 2025; Earth
@@ 2. Background & Motivation
2.1. Taint Bugs & Static Analysis
A taint bug is a type of vulnerability that arises when untrusted data is improperly handled within a program, leading to potential security risks. For example, in the Linux kernel, the ioctl system call often handles user input, without proper validation vulnerabilities can occur, such as buffer overflows and infinite loops (machiry_dr_2017, ; zhang_statically_2021, ).
We focus on taint bugs for several reasons. First, a wide variety of security vulnerabilities can be modeled as taint problems, including buffer overflows, integer overflows, use-after-free, and other memory corruption issues. Second, taint analysis provides a well-defined conceptual framework for reasoning about data flow from untrusted sources to sensitive operations (sinks). Third, the taint static analysis is both practical and widely adopted in security tools. These make taint analysis an ideal foundation for our work.
Suture (zhang_statically_2021, ) is a state-of-the-art static analysis tool designed specifically for detecting taint-style vulnerabilities in the Linux kernel. It employs a taint-tracking approach to identify potential vulnerabilities by analyzing the flow of untrusted data through the kernel code. Suture precisely tracks how data from user inputs, such as those passed via the ioctl system call, propagates through various functions and data structures, flagging any instances where this tainted data could lead to security issues based on their detectors.
Notably, Suture is an advanced taint analysis tool that can even analyze multi-entry data flows, which means it can track tainted data propagating through multiple system calls. Nevertheless, it still faces challenges in accurately determining whether the identified potential bugs are indeed exploitable, and still require considerable manual review to confirm their validity. We discuss these challenges in detail in the following parts of this section.
Figure 1. A potentially vulnerable code snippet from the Android kernel, simplified for illustration. The array values is tainted and all its elements are tainted as well.
[…]
This code snippet is a simplified version of a function from the Android kernel. It shows how the function retrieves the number of channels and their types, which could be exploited if not properly validated.
2.2. Motivating Example
The primary motivation for BugLens stems from the persistent challenge of achieving both precision and scalability in static vulnerability analysis, particularly for large-scale, complex systems like the Linux kernel. While static analysis is indispensable for identifying potential security issues, current tools often generate a high volume of findings, many of which are false positives (FPs) upon manual inspection. This significantly burdens security analysts and developers. BugLens aims to bridge this gap by introducing a methodology for more precise, context-aware evaluation of potential taint-style vulnerabilities, leveraging Large Language Models (LLMs) in a guided manner.
Figure 1 shows a simplified code snippet from the Android kernel. The function is called with an array of tainted variables *values (Note: the address represented by values is not tainted, but its contents are, from a copy_from_user). Thus, retrieved by GET_NEXT, the variable nums and command_id are tainted as well. Static analysis may flag potential bugs for the following detectors:
A standard static taint analyzer might flag the following issues:
@@ 2.2. Motivating Example
    Tainted Loop Bound (TLB): At line 10, the loop iterates based on the tainted variable nums. A very large value for nums (e.g., close to INT_MAX) could lead to excessive iteration, potentially causing a Denial-of-Service (DoS) or creating conditions for other vulnerabilities.
@@
    Tainted Pointer Dereference (TPD): At lines 11, the GET_NEXT macro is called. Inside this macro (line 4), the tainted pointer values (represented as ptr) is dereferenced (*ptr++) after being incremented. Since the number of increments depends on the loop controlled by tainted nums, the pointer values in the iteration is tainted as well (an implicit taint propagation). Therefore, it could be reported as a tainted pointer dereference, which may potentially lead to an out-of-bounds or arbitrary memory access.
2.2.1. Ground Truth Analysis
However, a deeper analysis reveals a more nuanced reality:
@@ 2.2.1. Ground Truth Analysis
    The TLB is a potential true positive. If nums is extremely large, but the data in values remains within the bounds defined by MAX_PP_PARAMS_SZ for many iterations, the loop can run for a long time. If the read command_ids are invalid (triggering the default case and continue on line 18), the if (rc) check on line 20 is bypassed for those iterations. The loop only terminates when i reaches nums. An extremely large nums could thus cause excessive computation, leading to a DoS.
@@
    The TPD is a false positive. The potential out-of-bounds read is prevented by the explicit bounds check within the GET_NEXT macro itself (line 3). It compares the incremented pointer address against the calculated param_max_offset (derived from the known buffer size MAX_PP_PARAMS_SZ) before the dereference occurs (line 4). If the check fails, rc is set, and the dereference is skipped.
2.3. Challenges in Static Analysis
Our motivating example, involving the analysis of TLB and TPD (illustrated in Figure 1), highlights the specific challenges stemming from this precision gap, which BugLens aims to address.
2.3.1. Challenge 1: Simplified Vulnerability Modeling
A primary source of imprecision is the reliance on simplified vulnerability detection modeling. To maintain scalability and avoid missing potential bugs, static analyzers may use overly simplified rules. For instance, an analyzer might flag any tainted pointer in the loop bound conditions as a potential DoS vulnerability, without considering the properties of the loop, such as the TLB example. Similarly, flagging all arithmetic operations involving tainted variables as potentially overflowing ensures coverage but generates noise.
The difficulty lies in defining and implementing effective and comprehensive rules within traditional static analysis frameworks. Vulnerabilities manifest in diverse and subtle ways. Consider the TLB example: a tainted loop bound might lead to Denial-of-Service (DoS) directly through excessive iterations, e.g., while (i < nums) where nums is tainted and set to a large value. Alternatively, the DoS might also happen when the tainted data is used in loop control statement such as continue or break. Precisely encoding rules to capture the vast number of unsafe code patterns, subtle contextual factors, and the impact of complex control flow is extremely challenging. As a result, this reliance on simplified heuristics inevitably leads to false positives in static analysis for security vulnerability detection.
@@ 2.3.2. Challenge 2: Over-Approximation of Constraints
The second major source of imprecision is the over-approximation of path and data constraints. To avoid the exponential complexity of exploring all possible execution paths and data states, static analyzers make coarse assumptions about path feasibility and data flow. While essential for scalability, this over-approximation means the analysis often considers execution paths that are actually infeasible in practice.
Consequently, warnings are generated for potential vulnerabilities on these infeasible paths, significantly inflating the false positive count. For example, an analyzer might flag the TPD scenario as dangerous without determining if the tainted pointer has been properly checked with its legal bounds before dereferencing. This imprecision results in a high volume of false positives, and forces developers to manually trace complex paths and data constraints to determine feasibility, undermining the efficiency of automated analysis.
@@ 2.4. The Opportunity in LLMs
Despite their potential, LLMs are not inherently suited for rigorous program analysis tasks out-of-the-box. Their reasoning can be fragile when dealing with complex program dependencies, and they lack systematic constraint-solving mechanisms. Simply asking an LLM a high-level question like “Is this static analysis warning a true positive?” often yields unreliable results.
Specifically, LLMs (even the most advanced ones, see §5.4) may struggle with intricate control and data flow dependencies. For instance, when analyzing the TLB example (Figure 1), an LLM might observe sanity checks on the tainted data and incorrectly conclude the warning is a false positive:
”There are sanity checks for the bound value… if the check fails, an error code is set, and the function jumps to cleanup… Therefore, the tool’s warning appears to be a false positive.”
However, as shown in the example, the LLM might fail to recognize that a continue statement within the loop body bypasses these checks under certain conditions, potentially allowing the infinite loop to occur. This highlights a critical limitation: LLMs may overlook the precise implications of control flow structures and data dependencies, leading to flawed conclusions.
Therefore, effectively leveraging LLMs for refining static analysis requires more than simple prompting. It necessitates a structured approach that guides the LLM’s reasoning process, teaching it relevant program analysis concepts and focusing its attention on critical code features and dependencies. This challenge motivates the design of our BugLens framework, which aims to orchestrate LLM analysis for reliable post-refinement.
2.5. Our Approach
The challenges outlined in our motivation, namely the imprecision stemming from Simplified Vulnerability Modeling (C1) and Over-Approximation of Constraints (C2) in scalable static analysis, and Reasoning Hurdles for LLMs (C3) necessitate a new strategy to tackle these challenges. To address these, we introduce BugLens, a framework designed specifically as a post-refinement for results from traditional static analyzers.
BugLens’s primary goal is to enhance the precision of static analysis by filtering their imprecise analysis results. The framework comprises the following components, each targeting specific challenges:
@@ 2.5. Our Approach
    Security Impact Assessor (SecIA): Addressing Simplified Vulnerability Models (C1). Instead of relying solely on predefined patterns, it uses LLMs to analyze the potential security impact if tainted data identified by static analysis were completely controlled by an attacker. It then evaluates whether the tainted value could potentially lead to security vulnerabilities (e.g., memory corruption, DoS), as detailed in §3.2. By focusing on semantic consequences rather than simplified heuristics, SecIA provides a more precise assessment of security impact, overcoming C1.
@@
    Constraint Assessor (ConA): Addressing Over-Approximation of Constraints (C2). ConA performs a targeted analysis of the (semantically) relevant constraints (e.g., sanitizers) along the data flow path (§3.3). It first identifies and collects the relevant constraints, then assesses whether these constraints effectively prevent the triggering of the specific vulnerabilities. The evaluation helps filter out false positives due to the over-approximation of path and data constraints, which are inherent in scalable static analysis (C2).
@@
    Structured Analysis Guidance (SAG): Addressing Reasoning Hurdles for LLMs (C3). Simply prompting LLMs for analysis is unreliable due to their potential to overlook complex control and data dependencies. SAG is designed to mitigate this (C3). It provides structured analysis guidance to the LLMs performing ConA. Leveraging domain knowledge from program analysis, SAG employs carefully constructed prompts and few-shot examples. SAG instantiates key analysis principles, demonstrating how to systematically dissect code, trace dependencies, and evaluate conditions, especially in complex scenarios. This guidance steers the LLM towards a more rigorous and reliable analysis process, making the use of LLMs within BugLens effective.
3. Design
3.1. Overview and Approach
3.1.1. Problem Scope & Goal
In this paper, we focus on identifying and analyzing taint-style vulnerabilities in operating system kernels. These vulnerabilities arise when untrusted data flows from external sources (e.g., ioctl() calls) to security-critical operations. A security-critical operation could be memory corruption such as buffer overflow, infinite loop, or other dangerous operations that could compromise the kernel when manipulated with malicious input.
As mentioned in §2, the goal of BugLens is to facilitate the (scalable but imprecise) static analysis of taint-style vulnerabilities by leveraging LLMs in a more effective way.
3.1.2. Direct Prompt & Design Rationale
Our investigation reveals fundamental limitations in simply asking LLMs to evaluate static analysis results directly. When presented with potential vulnerabilities identified by static analysis tools, a naive approach would be to directly prompt the LLM with questions like “Is this code vulnerable?” while providing the static analysis report and relevant code context. We term this basic approach the Direct Prompt.
Our experiments (detailed in §5.4) demonstrate that such Direct Prompts lead to a concerning number of false negatives — cases where the LLM incorrectly classifies actual vulnerabilities as safe code. The Direct Prompt approach, while intuitive, proves unreliable for vulnerability detection.
We hypothesize that this limitation stems from the fundamental nature of current LLMs. Trained primarily on predicting the next token based on vast datasets, these models excel at recognizing common syntactic and semantic patterns. However, this strength can become a weakness in complex analysis tasks. LLMs may develop heuristics based on surface-level features (mccoy_embers_2024, ); for example, identifying the presence of a “sanity check” construct often correlates statistically with safe code in training data. Consequently, the model might classify code containing such a check as safe without performing the deeper reasoning required to determine if the check is actually effective under all relevant execution paths. Real-world vulnerabilities often exploit exactly these scenarios: checks that are bypassable, incomplete, or rendered ineffective by intricate control and data flows. The model’s tendency to rely on learned statistical correlations (prabhakar_deciphering_2024, ), potentially driven by attention mechanisms prioritizing frequent patterns, may cause it to overlook these critical dependencies.
To address these reasoning limitations, we propose Structured Analysis Guidance (SAG) - a structured approach for guiding LLMs toward more rigorous and effective vulnerability analysis:
  • $\bullet$
    Guided Stepwise Vulnerability Analysis: We design Constraint Assessor (ConA) as a four-stage process that mirrors static analysis process. This stepwise workflow guides the LLM through: (1) identifying vulnerability reachability conditions, (2) collecting relevant constraints along taint paths, (3) summarizing each constraint’s preconditions and postconditions, and (4) evaluating whether these constraints effectively prevent the vulnerability. This systematic approach helps overcome the LLM’s tendency toward shortcut reasoning. We detail this workflow in §3.3.
  • $\bullet$
    Guided Path Condition and Data Constraint Analysis: To address challenging analysis where implicit path conditions and data constraints often elude naive prompting, we provide the LLM with detailed guidance and few-shot examples, which explicitly direct the LLM through analyzing complex control flow constructs, distinguishing between different types of constraints, and properly evaluating their effectiveness. This in-context learning approach helps the LLM identify subtle vulnerabilities. We elaborate on these techniques in §3.4.
@@ 3.1.2. Direct Prompt & Design Rationale
Figure 2. Overview of BugLens, showing (1) Security Impact Assessor (SecIA) first assesses the security impact of the potential bugs identified by static analysis, and then (2) Constraint Assessor (ConA) assesses the data constraints and evaluates if the bug is feasible. ConA is guided by (3) Structured Analysis Guidance (SAG) to reason on the code more effectively, and can interact with the (4) Project Knowledge Agent (PKA) to get information about the codebase on-demand.
@@
3.1.3. Design Components
Besides prompting LLMs with SAG, the rest of the BugLens framework consists of three components, as shown in Figure 2:
@@ 3.1.3. Design Components
    Security Impact Assessor (SecIA): This component evaluates the security impact of the potential bugs identified by static analysis. It identifies the Critical Operations (COp) that are influenced by the tainted data and filters out benign patterns.
@@
Additionally, BugLens also adopts some commonly-used prompting techniques, such as (1) Majority-vote querying, where we query the model multiple times and take the most common answer; (2) Chain-of-Thought (CoT) prompting (wei_chain–thought_2023, ); and (3) Schema-constrained summarization, where a follow-up prompt requests the model’s own output in a fixed XML format, making LLM’s response easy to parse for subsequent steps.

### v3

@@ Abstract.
Static analysis plays a crucial role in software vulnerability detection, yet faces a persistent precision-scalability trade-off. In large codebases like the Linux kernel, traditional static analysis tools often generate excessive false positives due to simplified vulnerability modeling and over-approximation of path and data constraints. While Large Language Models (LLMs) demonstrate promising code understanding capabilities, their direct application to program analysis remains unreliable due to inherent reasoning limitations.
We introduce BugLens, a post-refinement framework that significantly enhances static analysis precision for bug detection. BugLens guides LLMs through structured reasoning steps to assess security impact and validate constraints from the source code. When evaluated on Linux kernel’s taint-style bugs detected by static analysis tools, BugLens improves precision approximately 7-fold (from 0.10 to 0.72), substantially reducing false positives while uncovering four previously unreported vulnerabilities. Our results demonstrate that a well-structured, fully-automated LLM-based workflow can effectively complement and enhance traditional static analysis techniques.
conference: ; May 2025; Earth
@@ 1. Introduction
  • $\bullet$
    Open Source. We open-source our code and data of BugLens on GitHub [^1] https://github.com/seclab-ucr/BugLens.
@@ 2. Background & Motivation
2.1. Taint Bugs
Taint-style bugs involve insecure data propagation, where unsanitized data (taint source) reaches a sensitive program location (taint sink), causing flaws like out-of-bound access. For instance, an unchecked input integer (source) might become an array index (sink). Proper sanitization (e.g., range checks) is typically missing or insufficient. Such bugs can lead to severe vulnerabilities like buffer overflows or denial-of-service.
In the kernel, these bugs often occur when untrusted user input (e.g., syscall arguments) propagates to sensitive kernel operations (e.g., arithmetic calculations) without adequate sanitization (machiry_dr_2017, ; zhang_statically_2021, ). Detecting these bugs is challenging due to:
@@ 2.1. Taint Bugs
    Diverse and General Sinks. Unlike user-space applications with often well-defined sink APIs (e.g., exec()), the kernel’s high privilege means almost any operation (even a single arithmetic calculation) could be a sink if affected by unsanitized user data. This broad scope increases false positives. Additionally, low-level coding optimizations common in the kernel (e.g., intentional integer overflows) can complicate accurate detection.
@@
    Scattered and Intricate Sanitization. Sanitization checks can be distant from sinks, often crossing function boundaries within the large and complex kernel codebase. These checks are frequently intertwined with domain-specific kernel logic (e.g., privilege systems, configurations), posing significant hurdles for taint analysis tools.
2.2. Static Analysis for Taint Bugs
Suture (zhang_statically_2021, ) is the state-of-the-art work from academia aiming at kernel taint-style bug detection, which extends the previous work, Dr. Checker (machiry_dr_2017, ), by adding cross-entry taint tracking capability and improving analysis precision. CodeQL (codeql2025, ) is one of the most powerful and widely used industry code analysis engines, capable of taint analysis (taint_tracking_CodeQL, ) and query-based bug detection based on customizable rules. CodeQL has a large community that develops and maintains many different bug detection rules, including those for kernel taint-style bugs (e.g., (backhouse_stack_2018, )). Despite their effectiveness, all these tools exhibit high false alarm rate (e.g., ~90%), underscoring the inherent difficulties in accurately identifying kernel taint-style vulnerabilities. We elaborate on these difficulties and challenges in §3.
3. Challenges and Design Rationale
3.1. Challenges in Static Analysis
3.1.1. Challenge 1 (C1): Simplified Vulnerability Modeling
The first source of imprecision is the reliance on simplified vulnerability detection modeling. To maintain scalability and avoid missing potential bugs, static analyzers often employ overly simplified detection rules.
For instance, a static analyzer typically flags specific code patterns as potential bugs. A Tainted Arithmetic Detector (TAD) would identify any arithmetic operation involving tainted variables (e.g., var += size, where var is tainted) as a potential integer overflow vulnerability. However, in Linux kernel practice, these patterns frequently do not translate to exploitable vulnerabilities for two key reasons:
@@ 3.1.1. Challenge 1 (C1): Simplified Vulnerability Modeling
    The Behavior Itself is Benign. The kernel’s reliance on low-level C idioms, pointer manipulation, and intentional ‘unsafe’ design patterns often confounds traditional analyzers, leading to inaccuracies, e.g., our evaluation shows that 61.2% of the flagged ‘bugs’ were actually benign (§6.4.2). For example, in Linux drivers, there are cases where integer overflow is expected and benign, such as with jiffies:
    For these cases, even when integer overflow is possible due to the tainted variable timeout, it is not a security vulnerability, as time_after is defined as #define time_after(a, b) (b - a) > 0). Such false positives commonly arise from intentional design patterns in the Linux kernel, including struct hacks (deliberate out-of-bounds access), type casting, unions, and data structure operations (e.g., ptr->next, container_of).
@@
    The Tainted Data is Properly Validated before reaching the sink. For example, the Taint in the Loop Bound (TLB) pattern is common in the Linux kernel, but many include proper validation. Consider this array traversal: [Uncaptioned image]
    Without the assert check, this loop could lead to an out-of-bounds access. However, the assert ensures that the loop index never exceeds the array bounds, making it safe. Static analyzers that does not account for such checks (e.g., path-insensitive ones) will misclassify this as a vulnerability, leading to false positives.
@@ 3.1.2. Challenge 2: Over-Approximation of Constraints
As mentioned above, the tainted data is often checked before the sink, and with a safe range. However, static analysis tools often over-approximate the path and data constraints to ensure scalability, as the kernel’s immense scale and continuous evolution make precise static analysis impractical, and the logic of the sanitization could often be distant or complex. Alternatively, a common approach in static analysis (infer_2015, ) is to directly prune paths with checks, regardless of their actual effects, which leads to false negatives. The existence of checks is not always effective in preventing the bug from happening, e.g., we discovered some unreported bugs due to a plausible but ineffective check (see §6.3). Consider the function clamp_num() shown in the following, which is designed to sanitize the tainted variable num:
The function clamp_num only sanitizes the tainted variable num effectively when the config is set. Therefore, a precise static analysis needs to perform (i) a range analysis on the tainted variable num, and (ii) a path-sensitive analysis on clamp_num() to determine path conditions.
@@ 3.2. The Opportunity in LLMs
Recent works(wen_automatically_2024, ; li_iris_2025, ) also show that LLMs can be used to refine static analysis results (C1 & C2) in a straightforward manner, i.e., to directly ask LLM with “Is this case from static analysis report vulnerable?” followed by the code snippet.
@@
Despite their potential, LLMs are not inherently suited for rigorous program analysis tasks out-of-the-box. Their reasoning can be fragile when dealing with complex program dependencies, and they lack systematic constraint-solving mechanisms. Our experiments (detailed in §6.4) demonstrate that such a simple prompting strategy leads to high false negatives — cases where the LLM incorrectly classifies actual vulnerabilities as safe code, especially against the presence of complex control flow and data constraints.
We observe that the false negatives stem from the fundamental limitation of existing LLMs when tasked with sophisticated, structured, and symbolic reasoning. Specifically, with simple prompting, LLMs often rely on surface-level spurious features (mccoy_embers_2024, ) to reason about the vulnerability. For example, the presence of a sanity check often correlates statistically with safe code (many practical static analyzers use such heuristics to prune paths). We observe the model classifies code containing such a check as safe without performing the deeper reasoning required to determine if the check is actually effective under the current execution paths. Real-world vulnerabilities often exploit exactly these scenarios: checks that are bypassable, incomplete, or rendered ineffective by intricate control and data flows. The model’s tendency to rely on spurious patterns without performing the actual reasoning has been shown ineffective in tracking critical dependencies and results in unreliable predictions in many different domains (prabhakar_deciphering_2024, ).
3.3. Design Rationale
To address challenges C1 through C3, we introduce BugLens, a fully automated, multi-stage LLM-based framework designed specifically as a post-refinement layer for static analyzers. The key design philosophy of BugLens is to scale the test-time compute of LLMs by allocating more tokens for precise reasoning of the intended program behaviors based on the model’s understanding of program semantics (snell2024scaling, ). Specifically, BugLens consists of the following specific key components, each targeting a specific challenge:
@@ 3.3. Design Rationale
    Security Impact Assessor (SecIA): Addressing Simplified Vulnerability Models (C1). Instead of relying solely on predefined patterns, BugLens leverages LLMs to analyze the potential security impact if tainted data identified by static analysis were completely controlled by an attacker. It then evaluates whether the tainted value could potentially lead to security vulnerabilities (e.g., memory corruption, denial of service (DoS)), as detailed in §4.1. By focusing on semantic consequences rather than simplified vulnerability modeling, SecIA provides a more precise assessment of security impact.
@@
    Constraint Assessor (ConA): Addressing Over-Approximation of Path & Data Constraints (C2). ConA uses LLMs to analyze whether data constraints in the code are sufficient to prevent potential vulnerabilities from being triggered. By tracing how tainted data is processed and constrained, ConA performs heuristic reasoning (e.g., to find sanitizations with names check_x) to evaluate the effectiveness of these safeguards. This approach is designed to handle the complexity and scalability challenges of systems like the Linux kernel, where traditional formal methods may lack precision or scalability.
@@
    Structured Analysis Guidance (SAG): Addressing Reasoning Hurdles for LLMs (C3). As described in §3.2.1, simply prompting LLMs for analysis often leads to spurious reasoning. To this end, SAG grounds LLMs’ reasoning in ConA with the scaffold that describes the typical program analysis workflow. At a high level, SAG employs specific prompts to describe key analysis principles and include in-context examples to demonstrate how to systematically dissect code, trace dependencies, and evaluate conditions, especially in complex scenarios. This guidance constrains the LLM towards a more rigorous and reliable analysis process. §4.3 elaborates on the design of SAG.
@@
Figure 1. Overview of BugLens, showing (A) Security Impact Assessor (SecIA) first assesses the security impact of the potential bugs identified by static analysis, and then (B) Constraint Assessor (ConA) assesses the data constraints and evaluates if the bug is feasible. ConA is guided by (C) Structured Analysis Guidance (SAG) to reason on the code more effectively, and can interact with the (D) Project Knowledge Agent (PKA) to get information about the codebase on-demand.
@@
4. Design
BugLens first takes the static analysis report as input, which identifies the context of the potential bugs and the tainted data flow, and then performs the following steps to evaluate the potential bugs, as shown in Figure 1:
@@ 4. Design
    Security Impact Assessor (SecIA): This component evaluates the security impact of the potential bugs identified by static analysis. It identifies the Critical Operations (COp) that are influenced by the tainted data and classifies them as either Normal Code or Requires Constraint Analysis.
@@
    Structured Analysis Guidance (SAG): This component provides the scaffold to constrain the LLM through pre-defined code reasoning procedures to help it understand the code and analyze the constraints effectively.
  • $\bullet$
@@
Additionally, BugLens also adopts the common prompting techniques to scale the test-time compute, such as (1) Majority-vote querying, where we query the model multiple times and take the most common answer; (2) Chain-of-Thought (CoT) prompting (wei_chain–thought_2023, ); and (3) Schema-constrained summarization, where a follow-up prompt requests the model’s own output in a predefined XML format, making LLM’s response easy to parse.